# Towards Source Mapping for Zero-Knowledge Smart Contracts: Design and Preliminary Evaluation

### v1

@@ Abstract.
conference: The First Workshop on EXPlainable and REliable Software Systems (EXPRESS 2025); June 2025; Trondheim, Norwayccs: Software and its engineering Compilersccs: Security and privacy Formal security modelsccs: Theory of computation Program analysis
@@ 1. Introduction
Source mapping refers to the process of correlating compiled bytecode instructions with their corresponding high-level source code locations. This mapping is essential for debugging, auditing, and symbolic execution, as it enables developers and auditors to trace execution behavior back to source code statements.
While source mapping is a standard feature in Solidity compilers such as solc, it is currently absent in zkSolc. This omission introduces practical challenges for developers working on zero-knowledge-based smart contracts. Specifically, the zkSolc compilation pipeline involves multiple transformation stages, including translation to LLVM IR and subsequent generation of zkEVM bytecode. These stages introduce non-linear transformations, such as instruction reordering, arithmetic simplification, and control flow restructuring, which break the direct correspondence between source code and bytecode offsets that traditional source mapping techniques rely on.
@@ 3.1. Blockchain Scaling and zk-Rollups
Unlike solc, which directly compiles Solidity source code to EVM bytecode, zkSolc introduces an intermediate compilation stage based on LLVM IR. This additional stage enables advanced optimizations tailored to zk-SNARK constraint generation but introduces substantial transformations that complicate traceability between source code and compiled output.
Figure 1. zkSolc compilation pipeline (Labs, 2025a). Yul is an intermediate language used in Solidity compilation. Vyper and LLL are alternative smart contract languages supported by EVM toolchains but are outside the scope of this work.
@@ 3.3. Source Mapping in Solidity Compilers
Source mapping is a standard feature in traditional smart contract compilers, providing a correspondence between compiled bytecode instructions and their source-level origins. In conventional Solidity compilers such as solc, source maps are generated to facilitate debugging, symbolic execution, and security auditing. These maps enable developers to trace execution behavior back to Solidity source lines, supporting tools such as Remix IDE (Project, 2025b), Hardhat Debugger (hardhat), and Truffle (Suite, 2025), as well as formal analysis engines (of Bits, 2025c; ConsenSys, 2025).
@@ 3.4. Challenges in Source Mapping for Zero-Knowledge Compilation
These transformations disrupt the direct relationship between Solidity code and compiled zkEVM bytecode, rendering conventional offset-based source mapping techniques ineffective.
@@ 4. Methodology
This section presents the design of a preliminary source mapping framework for zkSolc. The framework aims to enable reliable traceability between Solidity source code, LLVM IR, and zkEVM bytecode, addressing the traceability challenges introduced by zero-knowledge compilation pipelines. Our approach is intended to facilitate debugging, security auditing, and performance analysis in zk-Rollup environments.
@@ 4.2. Mapping Algorithm and Data Structures
The mapping process is designed to track transformations at each compilation stage, preserving traceability despite the optimizations introduced by LLVM and zk-SNARK-specific rewrites.
@@ 4.3. Integration into zkSolc
To enhance usability, a runtime query API is introduced, allowing developers to retrieve source mappings dynamically. This API is designed to be compatible with existing debugging environments such as Remix and Hardhat, enabling runtime errors to be traced back to Solidity source code.
@@ 4.4. Illustrative Example
The mapping information is stored in a structured table:
@@
This mapping enables developers and auditors to correlate runtime execution behavior with Solidity source code and intermediate representations.
@@ 5.5. Alignment Criteria and Example
In cases where source constructs were eliminated or transformed during LLVM optimizations (e.g., constant folding or dead code elimination), no mapping entry was recorded.
@@ 6. Limitations
While the proposed source mapping framework for zkSolc demonstrates promising preliminary results, several limitations remain. These limitations highlight areas for further refinement and inform directions for future work.
One primary limitation is the compilation overhead introduced by the mapping process. Although our evaluation shows that the overhead remains moderate and acceptable for most development scenarios, it may pose challenges in high-throughput environments or automated deployment workflows involving frequent contract compilation. Further optimization of the mapping algorithm, particularly during the Solidity-to-LLVM IR translation phase, may help mitigate this overhead.
The current framework provides reliable traceability for standard Solidity constructs but exhibits reduced mapping accuracy when handling contracts with complex language features. Specifically, contracts involving deeply nested function calls, inline assembly, unconventional control flow, or experimental Solidity features present challenges for accurate traceability. Extending support for these advanced constructs without incurring significant performance penalties remains an open research problem.
Another practical limitation concerns integration with existing developer toolchains. Although the framework is designed to support execution trace reconstruction, it has not yet been fully integrated with widely used environments such as Remix, Hardhat, or Truffle. This limits its accessibility for developers accustomed to these platforms. Developing dedicated plugins or extensions to facilitate integration would substantially improve usability.
Furthermore, the current evaluation focuses exclusively on zkSolc and zk-Rollup-based compilation pipelines. The applicability of the framework to other zero-knowledge proof systems—such as Linea, Starknet, or Polygon zkEVM—has not been explored. While the underlying techniques may generalize to these platforms, confirming this requires further investigation and adaptation to system-specific compilation pipelines.
The framework’s mapping accuracy may also degrade in scenarios involving complex cross-contract interactions or deeply nested modifiers. Our evaluation identified minor inaccuracies in such cases, primarily due to compiler optimizations and non-linear control flow transformations. Improving mapping robustness under these conditions is an avenue for future enhancement.
Additionally, the current implementation focuses exclusively on static analysis and compilation-time mapping. It does not support runtime debugging features such as dynamic execution tracing or real-time error reporting. Incorporating such capabilities could significantly improve developer experience and is a promising direction for future work.
We also acknowledge that the evaluation datasets, although representative, may not fully capture all edge cases encountered in production environments. Expanding the evaluation to include a broader variety of contract types, code patterns, and execution scenarios would improve confidence in the framework’s generalizability.
Finally, the framework is tightly coupled with the current zkSolc architecture. Substantial changes to the compiler’s optimization strategies, IR transformations, or code generation processes may require corresponding modifications to the mapping framework to maintain accuracy and consistency.
Addressing these limitations will be the focus of future work. In particular, we plan to enhance mapping precision, reduce performance overhead, improve toolchain integration, and investigate generalization to other zero-knowledge compilation pipelines and execution environments.
@@ 7. Related Work
This section reviews prior research related to source mapping, smart contract debugging, and zero-knowledge Rollup compilation pipelines. While these areas have been studied independently, source mapping challenges specific to zero-knowledge-oriented compilers such as zkSolc remain underexplored.
@@ 7.1. Source Mapping and Debugging in Solidity
In Ethereum smart contract development, source mapping is a standard feature supported by mainstream Solidity compilers. For example, solc generates source maps that link compiled EVM bytecode instructions to corresponding Solidity source locations. This capability is fundamental to development environments and debugging tools such as Remix, Hardhat, and Truffle (Project, 2025b; Suite, 2025). Source mapping also plays a key role in formal verification workflows, enabling symbolic execution, execution trace reconstruction, and property checking.
Several analysis frameworks have been proposed to improve smart contract reliability and security. Static analyzers such as Slither (of Bits, 2025c), SmartCheck (SmartDec, 2025), and MythX (ConsenSys, 2025) detect common vulnerabilities at the source code level. Dynamic analysis tools, including Manticore (of Bits, 2025b) and Echidna (of Bits, 2025a), employ fuzzing and symbolic execution to expose runtime vulnerabilities. These tools fundamentally rely on accurate source mappings to correlate low-level execution behavior with high-level Solidity constructs. In compilation pipelines lacking source mapping support, such as zkSolc, debugging and auditing workflows are significantly hindered.
@@ 7.2. Zero-Knowledge Rollups and Compilation Tooling
Zero-knowledge Rollup (zk-Rollup) technologies have emerged as a leading solution for blockchain scalability (Ethereum Foundation, 2024). zk-Rollup systems, including zkSync (Labs, 2025b), process transactions off-chain and submit succinct cryptographic proofs to Ethereum for verification. Specialized compilers such as zkSolc have been developed to produce zero-knowledge-compatible bytecode tailored for these systems. Prior research in this space has predominantly focused on protocol-level efficiency, proof generation, and cryptographic correctness (Hacken, 2023). However, compiler-level challenges—particularly those concerning developer tooling and source-level traceability—have received comparatively limited attention.
A distinguishing characteristic of zero-knowledge compilers is their reliance on intermediate representations such as LLVM IR to facilitate aggressive optimization and proof-friendly transformations. These transformations introduce non-linear changes to control and data flow, complicating traceability between source code and compiled artifacts.
@@ 7.3. Source Mapping in Intermediate Representations
Source mapping between high-level languages and intermediate representations has been extensively studied in traditional compiler pipelines. For example, the LLVM framework (Lattner and Adve, 2004) supports debug metadata and source location annotations to enable traceability during optimization and code generation (Project, 2025a). Several studies have proposed techniques to preserve source mapping fidelity in the presence of conventional compiler optimizations, such as inlining and loop unrolling.
However, these techniques are not directly applicable to zero-knowledge compilation pipelines. The introduction of circuit-specific rewrites, arithmetic simplification at the bit level, and constraint generation for zk-SNARK proofs introduces additional layers of complexity beyond those encountered in general-purpose compilation.
@@ 7.4. Positioning and Contribution
Recent discussions within the zero-knowledge proof and smart contract development communities have emphasized the lack of robust debugging and traceability support in zk-Rollup ecosystems. To the best of our knowledge, no prior work has systematically addressed the problem of source mapping in the context of zero-knowledge-specific compilers such as zkSolc. Existing source mapping techniques in EVM or LLVM-based pipelines do not account for the non-linear transformations and circuit-level constraints inherent to zero-knowledge proof systems.
This paper takes an initial step toward addressing this gap. Our framework introduces a mapping mechanism tailored to the multi-stage compilation pipeline of zkSolc, providing preliminary support for tracing execution behavior across Solidity source code, LLVM IR, and zkEVM bytecode. While the approach does not offer formal correctness guarantees, it incorporates lightweight syntactic and structural validation to improve mapping consistency. We position this work as a foundation for future research toward more robust debugging, auditing, and verification toolchains in zero-knowledge smart contract development.
@@ 8. Conclusion
This paper presented a preliminary investigation into source mapping support for zkSolc, the compiler used to generate zero-knowledge-compatible smart contract bytecode. We proposed a source mapping framework that establishes traceability between Solidity source code, LLVM IR, and zkEVM bytecode, addressing a critical tooling gap in the zk-Rollup compilation pipeline.
Our evaluation on both benchmark and real-world Solidity contracts demonstrated that the proposed framework achieves high mapping accuracy while introducing only moderate compilation overhead. These results suggest that, despite the non-linear transformations inherent in zero-knowledge compilation workflows, meaningful source-level traceability can be recovered to support debugging, auditing, and performance analysis.
We acknowledge that this work constitutes an initial step toward robust source mapping in zero-knowledge environments. Several limitations remain, including reduced mapping precision in contracts featuring complex control flow and inline assembly, as well as integration challenges with existing developer toolchains. Furthermore, while we introduced lightweight syntactic and structural validation, formal guarantees of mapping correctness remain an open problem.
Future work will focus on addressing these limitations. Specifically, we plan to refine the mapping algorithm to improve accuracy for advanced Solidity constructs, optimize compilation overhead, and integrate the framework with popular development environments such as Remix and Hardhat. Extending the approach to other zero-knowledge compilation pipelines—such as those used by Linea, Starknet, and Polygon zkEVM—also represents an important direction for future research.
We hope that this preliminary investigation will lay the groundwork for further research on debugging, verification, and developer tooling in zero-knowledge smart contract ecosystems, ultimately contributing to a more transparent and reliable Layer 2 development experience.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: 34th ACM SIGSOFT International Symposium on Software Testing and Analysis; June 25–28, 2025; Trondheim, Norwaybooktitle: 34th ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA Companion ’25), June 25–28, 2025, Trondheim, Norwaydoi: 10.1145/3713081.3732932isbn: 979-8-4007-1474-0/2025/06ccs: Software and its engineering Compilersccs: Security and privacy Formal security modelsccs: Theory of computation Program analysis
@@ 1. Introduction
Source mapping refers to the process of correlating compiled bytecode instructions with their corresponding high-level source code locations. This mapping is essential for debugging, auditing, and symbolic execution, as it enables developers and auditors to trace execution behavior back to source code statements (Documentation, 2025).
While source mapping is a standard feature in Solidity compilers such as solc, it is currently absent in zkSolc (Labs, 2024). This omission introduces practical challenges for developers working on zero-knowledge-based smart contracts. Specifically, the zkSolc compilation pipeline involves multiple transformation stages, including translation to LLVM IR and subsequent generation of zkEVM bytecode. These stages introduce non-linear transformations, such as instruction reordering, arithmetic simplification, and control flow restructuring, which break the direct correspondence between source code and bytecode offsets that traditional source mapping techniques rely on.
@@ 2.3. Challenges in Implementing Source Mapping in zkSolc
Here, tmp2 is a temporary variable representing an intermediate value generated during IR-level translation of a Solidity expression. These transformations improve proof efficiency but obscure the original source semantics, making conventional source mapping infeasible.
@@ 3.1. Blockchain Scaling and zk-Rollups
Unlike solc, which directly compiles Solidity source code to EVM bytecode, zkSolc introduces an intermediate compilation stage based on LLVM IR (Labs, 2025a; Immutable, 2022). This additional stage enables advanced optimizations tailored to zk-SNARK constraint generation but introduces substantial transformations that complicate traceability between source code and compiled output.
Figure 1. zkSolc compilation pipeline (Labs, 2025b). Yul is an intermediate language used in Solidity compilation. Vyper and LLL are alternative smart contract languages supported by EVM toolchains but are outside the scope of this work.
@@ 3.3. Source Mapping in Solidity Compilers
Source mapping is a standard feature in traditional smart contract compilers, providing a correspondence between compiled bytecode instructions and their source-level origins. In conventional Solidity compilers such as solc, source maps are generated to facilitate debugging, symbolic execution, and security auditing. These maps enable developers to trace execution behavior back to Solidity source lines, supporting tools such as Remix IDE (Project, 2025b), Hardhat Debugger (Foundation, 2025), and Truffle (Suite, 2025), as well as formal analysis engines (of Bits, 2025c; ConsenSys, 2025).
@@ 3.4. Challenges in Source Mapping for Zero-Knowledge Compilation
These transformations disrupt the direct relationship between Solidity code and compiled zkEVM bytecode, rendering conventional offset-based source mapping techniques ineffective (Xiao et al., 2025).
@@ 4. Methodology
This section presents the design of a preliminary source mapping framework for zkSolc. The framework aims to enable reliable traceability between Solidity source code, LLVM IR, and zkEVM bytecode, addressing the traceability challenges introduced by zero-knowledge compilation pipelines. Our approach is intended to facilitate debugging, security auditing, and performance analysis in zk-Rollup environments (Fredrikson et al., 2014; Xiao et al., 2025).
@@ 4.2. Mapping Algorithm and Data Structures
The mapping process is designed to track transformations at each compilation stage, preserving traceability despite the optimizations introduced by LLVM and zk-SNARK-specific rewrites. Prior work on certifying compilers and domain-specific languages for secure proofs motivates preserving semantic traceability across compilation layers (Chiesa and Tromer, 2010; Bogdanov et al., 2022).
@@ 4.3. Integration into zkSolc
To enhance usability, a runtime query API is introduced, allowing developers to retrieve source mappings dynamically. This API is designed to be compatible with existing debugging environments such as Remix and Hardhat, enabling runtime errors to be traced back to Solidity source code. Efforts to standardize compiler output for zero-knowledge protocols highlight the importance of runtime traceability for debugging and auditing (Mccall, 2013; Shoup, 2010).
@@ 4.4. Illustrative Example
The mapping information is stored in a structured table that enables developers and auditors to correlate runtime execution behavior with Solidity source code and intermediate representations. Table 1 presents illustrative entries from the mapping table for the ZKVoting contract. Each row demonstrates how a high-level Solidity statement is translated through LLVM IR into zkEVM bytecode. The first column identifies the source-level construct, while the second shows its corresponding LLVM IR instruction. The third column indicates the bytecode offset at which the translated instruction appears, and the fourth records additional metadata relevant to zero-knowledge semantics, such as constraint checks. This table exemplifies the traceability our framework provides—enabling developers to understand how source-level logic is transformed and embedded in the final zkEVM output. By aligning semantic information across stages, it supports runtime debugging and static auditing in zkRollup-based systems.
@@ 4.5. Preliminary Consistency Validation
4.6. Handling Zero-Knowledge Compiler Transformations
To accurately map between Solidity source code and zkEVM bytecode via LLVM IR, our framework must address challenges that go beyond traditional AST-based techniques. Zero-knowledge-friendly compilation introduces several non-trivial transformations that can significantly alter control and data flow.
Instruction Reordering.
zkSolc applies transformation passes that reorder instructions to optimize circuit layout or constraint generation. Our mapping algorithm incorporates position tracking through metadata propagation and symbolic tagging across the LLVM IR layer. These tags help reconstruct the original source-level execution order despite backend reordering (Fredrikson et al., 2014).
Arithmetic Simplifications.
zkSolc aggressively applies simplifications such as constant folding and algebraic identity reduction to minimize proof size. We handle these cases by analyzing the def-use chains in LLVM IR and comparing them with the original Solidity expressions. If simplifications are detected, fallback annotations are used to indicate approximate rather than exact line mappings.
Control Flow Restructuring.
zkSolc may convert structured control flows (e.g., if-else, for) into flattened jump-based forms. We reconstruct the high-level structure by tracking basic block connectivity and dominator relationships within the LLVM IR control flow graph. Our mapping engine includes heuristics to remap these flattened constructs back to their corresponding high-level counterparts.
Zero-Knowledge Specific IR Passes.
Unlike traditional compilers, zkSolc includes constraint instrumentation passes. These insert intermediate variables and gadget calls into the IR. Our method explicitly tracks such injected instructions and excludes them from source mapping unless the inserted logic corresponds to visible Solidity semantics (e.g., require conditions).
Together, these strategies enable our mapping framework to remain robust in the face of proof-oriented rewrites and maintain traceability across the entire compilation pipeline.
@@ 5.5. Alignment Criteria and Example
In cases where source constructs were eliminated or transformed during LLVM optimizations (e.g., constant folding or dead code elimination), no mapping entry was recorded. Table 4 shows two representative examples from the ZKVoting contract that were manually validated for source mapping alignment. Each row lists a bytecode offset and the corresponding high-level Solidity statement. The third column indicates that these mappings were deemed correctly aligned, meaning the bytecode instruction at the given offset corresponds directly to the semantics and control flow of the specified source statement. For instance, the bytecode at 0x14 correctly reflects the execution of the require(verifyZKProof()) check, while the instruction at 0x20 corresponds to setting hasVoted[msg.sender] to true. This confirms that the mapping framework can accurately track and associate bytecode with its originating Solidity source under realistic contract execution.
@@ 6. Limitations
Our source mapping framework for zkSolc, while showing encouraging early results, still comes with several limitations that open up avenues for improvement and further exploration.
First, the additional compilation overhead introduced by the mapping process, although moderate in our benchmarks, could become a bottleneck in high-throughput or CI/CD-heavy environments. Optimizing the mapping pipeline—especially the translation from Solidity to LLVM IR—remains important for broader adoption in fast-paced development workflows.
Second, while the framework handles standard Solidity constructs with high accuracy, its performance declines in the presence of more complex patterns. Contracts that use deeply nested function calls, inline assembly, or less common control flows pose challenges for precise mapping. Supporting such features without compromising performance will require more refined static analysis techniques.
From a usability standpoint, integration with popular development tools like Remix, Hardhat, or Truffle is still lacking. Although our system supports execution trace reconstruction, the absence of plugins or dedicated interfaces means that developers working within these ecosystems must adopt additional tooling to benefit from our framework. Streamlining this integration will be key to making our solution developer-friendly.
Additionally, our current scope is limited to the zkSolc compiler and zk-Rollup settings. It remains unclear how well our approach generalizes to other zero-knowledge systems, such as Linea, Starknet, or Polygon zkEVM. Adapting to different compilation pipelines and proof systems may require rethinking parts of the mapping logic and warrants systematic evaluation.
We also observed minor inaccuracies when dealing with intricate contract structures, such as heavy use of modifiers or cross-contract calls. These issues are often tied to compiler optimizations and non-linear transformations, which can obscure traceability. Strengthening mapping robustness under such conditions is a challenge we’re keen to tackle.
Another current shortcoming is the lack of runtime debugging features. Our framework focuses entirely on static, compile-time mappings, and does not yet offer dynamic execution tracing or live error feedback. Bringing runtime visibility into the picture would significantly enhance the developer experience.
Furthermore, while our evaluation datasets are representative in scope, they may not cover all edge cases that arise in production-level contracts. A more comprehensive evaluation—including diverse contract types and usage scenarios—would help validate the framework’s robustness.
Lastly, because the implementation is tightly coupled with the internals of zkSolc, any significant updates to the compiler’s IR structure or optimization strategies may require parallel updates to our mapping logic. Maintaining compatibility in the face of upstream changes is a practical concern going forward.
We see these limitations not as drawbacks, but as opportunities for continued progress. Future work will focus on improving precision, reducing overhead, expanding tool support, and validating the approach across different platforms and real-world projects.
@@ 7. Related Work
We review relevant research spanning source mapping in Solidity, smart contract debugging, and compiler pipelines for zero-knowledge Rollups. While these domains have each attracted significant attention, source-level traceability in zero-knowledge-oriented compilers like zkSolc remains relatively unexplored.
@@ 7.1. Source Mapping and Debugging in Solidity
In Ethereum development, source mapping has become a foundational feature supported by the standard Solidity compiler solc, which emits source maps linking EVM bytecode back to Solidity source lines. This mapping is essential for development tools such as Remix, Hardhat, and Truffle (Project, 2025b; Suite, 2025), and underpins debugging, symbolic execution, and formal verification workflows.
A range of static and dynamic analysis tools have been proposed to improve smart contract reliability. Static analyzers such as Slither (of Bits, 2025c), SmartCheck (SmartDec, 2025), and MythX (ConsenSys, 2025) identify vulnerabilities at the source level, while dynamic tools like Manticore (of Bits, 2025b) and Echidna (of Bits, 2025a) rely on fuzzing and symbolic execution to surface runtime issues. All of these tools depend on accurate source mappings to connect low-level execution traces with high-level semantics. Recent work has expanded on semantic-based analysis (Grishchenko et al., 2018), scalable vulnerability detection (Tsankov et al., 2018; Jiang et al., 2018), Solidity security patterns (Wöhrer and Zdun, 2019), and improved debugging workflows through enhanced source mapping techniques (Zhang et al., 2021). In compiler pipelines that lack such mappings—as is currently the case with zkSolc—this connection breaks down, complicating debugging and auditing.
@@ 7.2. Zero-Knowledge Rollups and Compilation Tooling
Zero-knowledge Rollups (zk-Rollups) have emerged as a promising solution to blockchain scalability (Ethereum Foundation, 2024), with systems like zkSync (Labs, 2025c) offloading transaction processing and submitting succinct validity proofs on-chain. Recent surveys have emphasized the central role of zk-Rollups in scaling Ethereum (Xu et al., 2022), while research into zk-SNARK efficiency for smart contract verification highlights critical compiler optimization challenges (Chang et al., 2021). To support these systems, specialized compilers like zkSolc generate zero-knowledge-compatible bytecode. Prior research in this area has primarily focused on improving proof efficiency and cryptographic correctness (Hacken, 2023). In contrast, developer tooling and compiler support for source-level debugging have received limited attention.
Unlike standard compilers, zero-knowledge compilers often rely on intermediate representations such as LLVM IR to perform aggressive, proof-oriented optimizations. These optimizations reshape both control flow and data dependencies in non-linear ways, introducing new challenges for source-level traceability.
@@ 7.3. Source Mapping in Intermediate Representations
The problem of maintaining source mappings through compiler transformations has been extensively studied in traditional compilation settings. LLVM, for example, supports debug metadata and source annotations to track provenance through optimization stages (Lattner and Adve, 2004; Project, 2025a). Various techniques have been proposed to preserve mapping fidelity during inlining, loop unrolling, and other common transformations. Moreover, security verification at the compiler IR level remains an active research area (Wang et al., 2022).
However, such techniques do not directly translate to zero-knowledge compilers, which introduce additional layers of transformation such as arithmetic circuit rewrites, bit-level simplifications, and constraint encoding. These changes fundamentally alter the structure of the compiled program and obscure its connection to the original source code.
@@ 7.4. Positioning and Contribution
While source mapping is well understood in EVM and general-purpose LLVM contexts, its role in zero-knowledge systems has largely been overlooked. Discussions within the zk community have increasingly pointed to the lack of mature debugging and traceability support for zk-Rollup workflows. To our knowledge, no existing work systematically tackles the problem of source mapping within zero-knowledge compilers like zkSolc. The nonlinear and proof-centric nature of such compilers introduces challenges that conventional mapping strategies do not address.
7.5. Integration with Developer Toolchains
Looking ahead, we see strong potential in integrating source mapping capabilities into mainstream development tools. For instance, exporting mappings in Solidity-compatible formats (e.g., JSON source maps) could enable support in Remix, VSCode plugins, or Hardhat’s debugger. At a lower level, LLVM IR mappings and bytecode offsets could be surfaced via command-line tools for use in CI workflows or custom scripts within zkSync’s CLI environment. Such integrations would bridge the gap between advanced compilation infrastructure and practical developer needs.
@@ 8. Conclusion
In this paper, we explored the feasibility of supporting source-level traceability within zkSolc, the compiler behind zero-knowledge-compatible smart contracts. We introduced a source mapping framework that links Solidity code to both LLVM IR and zkEVM bytecode—addressing a notable gap in the zk-Rollup toolchain.
Through experiments on both curated benchmarks and real-world contracts, we showed that accurate source mappings can be recovered even in the presence of the aggressive and non-linear transformations typical of zero-knowledge compilation. The observed accuracy and reasonable compilation overhead suggest that meaningful debugging and auditing support is achievable for zk-based systems.
That said, our work represents an early step. Limitations remain in handling contracts with intricate control flows, heavy use of inline assembly, or uncommon language features. Our prototype also lacks full integration with widely used tools such as Remix or Hardhat, and we do not yet offer formal guarantees of correctness. These challenges point to rich areas for future refinement.
Looking ahead, we aim to enhance mapping precision, further reduce performance overhead, and build integrations with mainstream developer environments. We are also interested in extending our approach to other zk compilers such as those powering Starknet, Linea, and Polygon zkEVM, each of which introduces its own compilation idiosyncrasies.
Ultimately, we hope this work lays the foundation for more robust debugging, verification, and tooling support in zero-knowledge smart contract development. Our implementation will be made open source upon publication, enabling the community to build on, critique, and extend our results. We believe that a shared effort toward transparency and usability will accelerate the maturity of zk-Rollup ecosystems.